1. Demand and Supply

 


 

a.     In this market, what would be the equilibrium price and the equilibrium quantity? 2 Marks

 

b.    If price is $10, what would be the quantity demanded? 2 Marks

 

c.     If price is $25, what would be the quantity supplied? 2 Marks

 

d.    If the price is $25, would there be a surplus or a shortage? What would be the amount of the surplus or the shortage? 2 Marks


1
Expert's answer
2022-02-28T09:39:12-0500

Let be


Qs=-5+2p

Q d=30-5P

a.Qs=Qd

-5+2P=30-5P

7P=35

P=5

Q=5


b.

P=10

Qd=-20


c.

P=25

Qs=45


d.

 If the price is $25 would there be a surplus.

"surplus=25\times45=1125"